Can the surface x5y3 + x2y + z = 3 be expressed as the graph of a function x = f(y, z) near the point (2, 1/2, −3)? If yes, compute fy(1/2, −3) and fz(1/2, −3).


1
Expert's answer
2021-09-24T06:39:06-0400

Let "F(x,y,z)=x^5y^3 + x^2y + z- 3", and "P=(2,1\/2, -3)". Then

"dF(x,y,z)=(5x^4y^3 + 2xy)dx +(3x^5y^2 + x^2)dy + dz"

"dF(P)=12dx +28dy + dz"

Equation "dF(P)=12dx +28dy + dz=0" is the equation of the tangent plane to the surface F=0 at the point P.

Since "\frac{\partial F}{\partial x}(P)=12\ne 0" is invertible, then, by the implicit function theorem, the surface F=0 locally, in a neiborhood of (yz)-projection of the point P, can be expressed as a graph of differentiable function "x=f(y,z)".

The tangent plane to the graph "x=f(y,z)" at the point "P'=(1\/2,-3)" has equation

"dx-\frac{\partial f}{\partial y}(P')dy-\frac{\partial f}{\partial z}(P')dz=0"

But we know that the tangent plane to the surface F=0 at the point P has equation

"12dx +28dy + dz=0"

We have two formulas of the same plane. Therefore, they must be proportional:

"{1}:{12}=-\frac{\partial f}{\partial y}(P'):28=-\frac{\partial f}{\partial z}(P'):1"

Therefore,

"\frac{\partial f}{\partial y}(P')=-28/12=-7/3"

"\frac{\partial f}{\partial z}(P')=-1/12"